2. (6 pts) There are many different types of phages as they can RNA or DNA, and may be single stranded or double straned


A) (3 pts) You isolate an unknown phage and fine it to have 20% A, 32% T, 25% G, and 33% C. What does this tell you regarding the classification of this phage? Why?

           
B) (3 pts) This phage infects bacteria by integrating its DNA into the bacterial chromosome. The rate of success of this integration is dependent on the topology of the bacterial DNA, and this rate is high (> 70%) when the superhelical density, σ, of the host DNA is less than -0.06, but poorly integrates when σ is greater than -0.045. Plasmid DNA isolated from the bacterial culture is found to be 13,800 bp and have an Lk of 1,222. Predict the likelihood that this phage will be able to infect the bacteria? Why?
Answer #1

A) The phage belongs to single stranded DNA phage. This is because the percentage of A is not equal to percentage of C similarly percentage of C is not equal to percentage of G. If a DNA molecule double stranded then it should have the same percentage of purines wit their corresponding pyramidines.

B) Given that,

Base pair= 13800 bp

Lk= 1222

Lko=base pair/10.4

Lko=13800/10.4 = 1326.92

Change in linking no. \Delta LK= LK-LKo

\Delta LK= 1222-1326.92= -104.92

Now,

Superhelical density \sigma = \Delta LK/LKo= -104.92/1326.92= -0.079

The superhelical density less than -0.06, so the rate of integration is more than 70% and the chances of infecting the bcteria is high.